ericcrawford:

A sled travels 15 meters down a slope inclined at 30° with the horizontal. What is the horizontal displacement of the sled?

Mercury:

 cos(theta) = adjacent/hypotenuse

ericcrawford:

uhhhhhh

Mercury:

theta is the angle x is the adjacent side (which is what you want to solve for, the horizontal displacement) 15 is the hypotenuse solve for x

ericcrawford:

I'm still confused

Mercury:



Mercury:

cos(30) = x/15 solve for x

ericcrawford:

2?

Mercury:

cos(30) = x/15 multiply both sides by 15 to get x = cos(30) * 15 = ?

ericcrawford:

30?

Mercury:

remember that cos(30) is sqrt(3)/2

Mercury:

x = cos(30) * 15 = sqrt(3)/2 * 15 or about 12.99 in decimal form

Mercury:

no, I just said x = sqrt(3)/2 * 15 or about 12.99 in decimal form

ericcrawford:

oh look I am awful at algebra

ericcrawford:

thank you